iii) 4/x + 3y = 14

3/x -4y = 23

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Putting in the given equation we get, So, 4m + 3y = 14 => 4m + 3y – 14 = 0 ……………..…..(1) 3m – 4y = 23 => 3m – 4y – 23 = 0 ……………………….(2) By cross-multiplication, we get, m/(-69-56) = y/(-42-(-92)) = 1/(-16-9) -m/125 = y/50 = -1/ 25 -m/125 = -1/25 and y/50 = -1/25 m = 5 and b = -2 m = 1/x = 5 So , x = 1/5 y = -2